The average repair cost of a random sample of 25 refrigerators is $150 with a standard deviation $15. What is the margin of error of the 90% confidence interval for the population mean cost constructed by using this sample? Enter a dollar amount, round to the nearest cents
Answer #1

Solution no 25 Mean (2): 150 Standard deviation (2):15 At 90% confidence level a: 1-907-1-0.90 .0.1 d2=0.05 df anal=25-1 =24